OpenStudy (anonymous):

Volume of the object = 192in^3 What is the equation for the volume of the object written in terms of x ?

OpenStudy (anonymous):

http://www.math.com/tables/geometry/volumes.htm Are you just looking for the formula?

OpenStudy (anonymous):

no not really, just the equation for the volume of the object written in terms of x. L = 8 W = 2 H = 12

OpenStudy (anonymous):

just post what you think thanks! :D brb

OpenStudy (anonymous):

I got nothing

OpenStudy (anonymous):

dude I thought u said this was LGB's question lol XD

OpenStudy (anonymous):

liar

OpenStudy (anonymous):

I asked LGBA to help with this question..slacker

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

>xD

OpenStudy (anonymous):

your eyes are cut off in the pic AWW!!!!!!!!!

OpenStudy (anonymous):

a mad laughing face?

OpenStudy (anonymous):

WTF is >xD?

OpenStudy (anonymous):

ur don;t lie

OpenStudy (anonymous):

the mad laughing face ...we can't chat on a question though so msg me

OpenStudy (anonymous):

geez how can be at some many places at once? msg, chat now here? lol

OpenStudy (anonymous):

uh..

OpenStudy (anonymous):

I msgd lbasallote he should be able to help me bc I don't think I quite understand what you're asking

OpenStudy (anonymous):

I myself don't understand what I'm asking, since this is a question from my Algebra lesson.

OpenStudy (anonymous):

haha touché

OpenStudy (anonymous):

I'm not asking nothing

OpenStudy (anonymous):

LGB :D

OpenStudy (anonymous):

BECCA :D

OpenStudy (anonymous):

TRIBUTIZED :D

OpenStudy (anonymous):

wut?

OpenStudy (lgbasallote):

i think it woiuld be x = (8)(2)(12) x = 192 in^3 ? \( \color{green}{\mathsf{L}}\) \(\small \color{red}{\mathcal{O}}\) \(\color{blue}{\mathsf{L}}\)

OpenStudy (anonymous):

^that was gonna be my second guess

OpenStudy (anonymous):

i think it's x = l * w * h

OpenStudy (anonymous):

so.. if the volume is 192 and l = 8 w = 2 h = 12 then.. i think it's going to be 192 = 8*2*12

OpenStudy (anonymous):

Correct :-)

OpenStudy (anonymous):

oh goodie! thanks!